OpenStudy (anonymous):

There are at most 12 bicycles and tricycle in a school playground. There are at least 17 wheels altogether. Let b equal the number of bicycles and t equals the number of tricycle. Which system describes this situation? Why?

OpenStudy (michele_laino):

please, keep in mind that foreach bicycle we have two wheels, and for each ticycle we have thre wheels furthermore fm your data,we can write: b+t=17

OpenStudy (michele_laino):

oops sorry... b+t=12

OpenStudy (michele_laino):

please, yoiu have to add another relation between b and t

OpenStudy (michele_laino):

number of wheels due to bicycle are 2*, whereas numbers of wheelsdue to tricycle ae 3*t, so?

OpenStudy (michele_laino):

2*b+3*t=...

OpenStudy (michele_laino):

please write the right side of the above equation

OpenStudy (michele_laino):

oops sorry number of wheels due to bicycle is 2*b..

OpenStudy (michele_laino):

please complete the equation: 2*b+3*t=...